OpenStudy (anonymous):

Please!! Describe how to find the percent of change in the price of a skateboard from last year to this year.

hero (hero):

In general, to find percent change, use the following formula: \[\% \Delta = \frac{|a - b|}{a} \times 100\] a = original amount b = new amount

hero (hero):

Sorry it took so long, equation editor malfunction

OpenStudy (anonymous):

it's ok

hero (hero):

You're welcome